OpenStudy (jacobbirddog):

Plz Help! ill medal and fan you! The sum of two consecutive integers is 67. What are the numbers? A. 34 and 35 B. 33 and 34 C. 32 and 33 D. 32 and 35

OpenStudy (anonymous):

33 and 34

OpenStudy (anonymous):

Let the numbers be x,x+1 x+x+1=67 2x=67-1=33 x=33 numbers are 33,33+1 or 33,34